# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v1

@@ C.1 Stability of the projection map $(A,G)\mapsto-A^{-1}G$
$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\;\leq\;2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}\;+\;2\kappa^{2}\,\big(M+\|\widetilde{G}\|_{L^{\infty}(D)}\big)\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$ (97)
@@
Throughout, $\|\cdot\|$ denotes the operator norm induced by the Euclidean norm. For a matrix-valued function $M:D\to\mathbb{R}^{d\times d}$, write
$\|M\|_{L^{\infty}(D)}:=\operatorname*{ess\,sup}_{z\in D}\|M(z)\|,\hskip 18.49988pt\|M\|_{L^{2}(\mu)}:=\Big(\int_{D}\|M(z)\|^{2}\,\mu(dz)\Big)^{1/2},$
and similarly for vector-valued functions.
Step 1: Invertibility and inverse bound for $\widetilde{A}$. Fix $z\in D$ such that $A(z)$ is invertible (this holds for $\mu$-a.e. $z$). Let $E(z):=\widetilde{A}(z)-A(z)$. By (i) and (iii),
$\|A(z)^{-1}E(z)\|\leq\|A(z)^{-1}\|\,\|E(z)\|\leq\|A^{-1}\|_{L^{\infty}(D)}\,\|\widetilde{A}-A\|_{L^{\infty}(D)}\leq\kappa\cdot\frac{1}{2\kappa}=\frac{1}{2}.$
Hence $I+A(z)^{-1}E(z)$ is invertible and admits the Neumann-series inverse. Therefore,
$\widetilde{A}(z)^{-1}=(A(z)+E(z))^{-1}=(I+A(z)^{-1}E(z))^{-1}A(z)^{-1},$
and
$\|\widetilde{A}(z)^{-1}\|\leq\frac{1}{1-\|A(z)^{-1}E(z)\|}\,\|A(z)^{-1}\|\leq\frac{1}{1-1/2}\,\kappa=2\kappa.$
Taking the essential supremum over $z\in D$ yields
$\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\leq 2\kappa.$
Step 2: A pointwise bound for $\widetilde{A}^{-1}-A^{-1}$. For $\mu$-a.e. $z\in D$ where both inverses exist,
$\widetilde{A}(z)^{-1}-A(z)^{-1}=\widetilde{A}(z)^{-1}\big(A(z)-\widetilde{A}(z)\big)A(z)^{-1}.$
Thus,
$\|\widetilde{A}(z)^{-1}-A(z)^{-1}\|\leq\|\widetilde{A}(z)^{-1}\|\,\|\widetilde{A}(z)-A(z)\|\,\|A(z)^{-1}\|\leq(2\kappa)\,\|\widetilde{A}(z)-A(z)\|\,\kappa=2\kappa^{2}\|\widetilde{A}(z)-A(z)\|.$
Consequently,
$\|\widetilde{A}^{-1}-A^{-1}\|_{L^{2}(\mu)}\leq 2\kappa^{2}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$
Step 3: Control error bound. Recall $\pi=-A^{-1}G$ and $\widetilde{\pi}=-\widetilde{A}^{-1}\widetilde{G}$. Then
@@
Taking $L^{2}(\mu)$ norms and using Hölder ($L^{\infty}\times L^{2}\to L^{2}$) gives
@@
Using $\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\leq 2\kappa$ (Step 1), $\|G\|_{L^{\infty}(D)}\leq M$ (assumption (ii)), and Step 2, we obtain
$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\leq 2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}+2\kappa^{2}\,M\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$
Finally, since $M\leq M+\|\widetilde{G}\|_{L^{\infty}(D)}$, this implies (97). ∎
C.2 Slab-wise small-gain for the $q$-aggregated projection inputs
@@
In the portfolio problem the working domain $D$ carries a time coordinate; for concreteness, assume
$D\subset[0,T]\times\mathcal{S},\hskip 18.49988pt\mu(dt,d\xi)=dt\otimes\nu(d\xi),$
for some reference measure $\nu$ on $\mathcal{S}$. Fix a partition $0=t_{0}<t_{1}<\cdots<t_{K}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$ and define
$D_{k}:=D\cap\big([t_{k-1},t_{k}]\times\mathcal{S}\big),\hskip 18.49988pt\mu_{k}:=\mu|_{D_{k}},\hskip 18.49988pt\|f\|_{k}:=\|f\|_{L^{2}(\mu_{k})}.$
@@
Let $\mathcal{U}$ be a neighborhood of $\pi^{\star}$ in the deployable $\theta$-blind policy class such that for all $\pi\in\mathcal{U}$,
$\|A_{\pi}^{-1}\|_{L^{\infty}(D)}\leq\kappa,\hskip 18.49988pt\|G_{\pi}^{\mathrm{mix}}\|_{L^{\infty}(D)}\leq M_{G}.$
Assume that on each slab $D_{k}$ the $q$-aggregated projection inputs satisfy
$\|A_{\pi_{1}}-A_{\pi_{2}}\|_{k}\leq\bar{L}_{A}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988pt\|G_{\pi_{1}}^{\mathrm{mix}}-G_{\pi_{2}}^{\mathrm{mix}}\|_{k}\leq\bar{L}_{G}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},$ (98)
for all $\pi_{1},\pi_{2}\in\mathcal{U}$ and constants $\bar{L}_{A},\bar{L}_{G}>0$ that depend only on band data. Define
$\rho(\tau):=\Big(\kappa\bar{L}_{G}+\kappa^{2}M_{G}\bar{L}_{A}\Big)\tau^{1/2}.$ (99)
Then for each slab $D_{k}$ and all $\pi_{1},\pi_{2}\in\mathcal{U}$,
$\|T(\pi_{1})-T(\pi_{2})\|_{k}\;\leq\;\rho(\tau_{k})\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988ptT(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}.$ (100)
In particular, if the partition is chosen with $\max_{k}\tau_{k}\leq\tau^{\star}$ for some $\tau^{\star}>0$ such that $\rho(\tau^{\star})<1$, then
$\rho_{*}:=\max_{1\leq k\leq K}\rho(\tau_{k})<1$
and $T$ is a contraction on every slab with constant at most $\rho_{*}$.
@@
Fix $k$ and $\pi_{1},\pi_{2}\in\mathcal{U}$. Write $A_{i}:=A_{\pi_{i}}$ and $G_{i}:=G_{\pi_{i}}^{\mathrm{mix}}$. Then
$T(\pi_{1})-T(\pi_{2})=-A_{1}^{-1}(G_{1}-G_{2})-(A_{1}^{-1}-A_{2}^{-1})G_{2},$
and
$A_{1}^{-1}-A_{2}^{-1}=A_{1}^{-1}(A_{2}-A_{1})A_{2}^{-1}.$
Using Hölder ($L^{\infty}\times L^{2}\to L^{2}$) on $D_{k}$ together with $\|A_{i}^{-1}\|_{L^{\infty}(D)}\leq\kappa$ and $\|G_{2}\|_{L^{\infty}(D)}\leq M_{G}$, we obtain
@@
Applying (98) yields (100) with $\rho(\tau_{k})$ as in (99). ∎
Remark 3 (Verification of the $\tau^{1/2}$ gain and relation to prior slab analyses).
The $\tau^{1/2}$-gain in (98) is the same short-time parabolic smoothing effect used in our prior PGDPO analysis (see, e.g., huh2025breaking): one combines a Duhamel/semigroup representation of the relevant adjoint/costate objects with Young-type convolution bounds to obtain a factor $\tau^{1/2}$ on each short slab. In the present paper, the only additional bookkeeping is that $(A_{\pi},G_{\pi}^{\mathrm{mix}})$ are $q$-aggregated (in particular, linear expectations over $\theta$), which does not alter the semigroup estimates; it only changes constants through coefficient bounds uniform in $\theta$ on the compact parameter set.
C.3 Proof of Theorem 3
Setup.
Let $D$ be the working domain with reference measure $\mu$ and slab decomposition $\{D_{k},\mu_{k},\|\cdot\|_{k}\}_{k=1}^{K}$ as in Appendix C.2. For a deployable $\theta$-blind policy $\pi$, write $(A_{\pi},G_{\pi}^{\mathrm{mix}})$ for the $q$-aggregated projection inputs corresponding to the mixed-moment aggregation in (65)–(66). Define the projection map
$T(\pi)(z):=-A_{\pi}(z)^{-1}G_{\pi}^{\mathrm{mix}}(z),\hskip 18.49988ptz\in D,$
whenever $A_{\pi}(z)$ is invertible.
Let $\pi^{\mathrm{warm}}$ be the warm-up policy and set
$A_{\mathrm{warm}}:=A_{\pi^{\mathrm{warm}}},\hskip 18.49988ptG_{\mathrm{warm}}:=G_{\pi^{\mathrm{warm}}}^{\mathrm{mix}},\hskip 18.49988pt\pi_{\mathrm{proj}}:=T(\pi^{\mathrm{warm}})=-A_{\mathrm{warm}}^{-1}G_{\mathrm{warm}}.$
@@ C.3 Proof of Theorem 3
Let $\widehat{A}_{t}$ and $\widehat{G}_{t}^{\mathrm{mix}}$ be the BPTT/Monte Carlo estimators constructed under $\pi^{\mathrm{warm}}$, and denote
$\widehat{\pi}^{\mathrm{agg,mix}}:=-\widehat{A}_{t}^{-1}\widehat{G}_{t}^{\mathrm{mix}}\hskip 18.49988pt\text{on }D.$
Proof.
@@
Step 0 (Fixed-point form of the deployable optimum). Let $\pi^{\star}:=\pi^{\star,\mathrm{blind}}$ be a locally optimal interior deployable $\theta$-blind policy for the fixed-$q$ ex–ante problem. By the $q$-aggregated stationarity (Theorem 1), $\pi^{\star}$ satisfies
$A_{\pi^{\star}}(z)\,\pi^{\star}(z)=-G_{\pi^{\star}}^{\mathrm{mix}}(z)\qquad\text{for }\mu\text{-a.e.\ }z\in D,$
hence (under invertibility on $D$) it is a fixed point of $T$:
$\pi^{\star}(z)=T(\pi^{\star})(z)=-A_{\pi^{\star}}(z)^{-1}G_{\pi^{\star}}^{\mathrm{mix}}(z),\hskip 18.49988pt\mu\text{-a.e.\ }z\in D.$
Step 1 (Triangle decomposition). Add and subtract $\pi_{\mathrm{proj}}$:
$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star}\|_{L^{2}(\mu)}\leq\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi_{\mathrm{proj}}\|_{L^{2}(\mu)}+\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{L^{2}(\mu)}.$ (101)
Step 2 (Estimation error via Proposition 1). Apply Proposition 1 with
$(A,G)=(A_{\mathrm{warm}},G_{\mathrm{warm}}),\hskip 18.49988pt(\widetilde{A},\widetilde{G})=(\widehat{A}_{t},\widehat{G}_{t}^{\mathrm{mix}}).$
Assume the perturbative regime on $D$:
$\|A_{\mathrm{warm}}^{-1}\|_{L^{\infty}(D)}\leq\kappa,\qquad\|G_{\mathrm{warm}}\|_{L^{\infty}(D)}\leq M_{G},\qquad\|\widehat{A}_{t}-A_{\mathrm{warm}}\|_{L^{\infty}(D)}\leq(2\kappa)^{-1},\qquad\|\widehat{G}_{t}^{\mathrm{mix}}\|_{L^{\infty}(D)}\leq M_{G}.$ (102)
Then (97) yields
$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi_{\mathrm{proj}}\|_{L^{2}(\mu)}$ $\leq 2\kappa\,\|\widehat{G}_{t}^{\mathrm{mix}}-G_{\mathrm{warm}}\|_{L^{2}(\mu)}+4\kappa^{2}M_{G}\,\|\widehat{A}_{t}-A_{\mathrm{warm}}\|_{L^{2}(\mu)}.$ (103)
By the definition of $\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta})$ in Theorem 3,
$\|\widehat{A}_{t}-A_{\mathrm{warm}}\|_{L^{2}(\mu)}+\|\widehat{G}_{t}^{\mathrm{mix}}-G_{\mathrm{warm}}\|_{L^{2}(\mu)}\leq\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),$
hence
$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi_{\mathrm{proj}}\|_{L^{2}(\mu)}\leq C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),\hskip 18.49988ptC_{2}:=2\kappa+4\kappa^{2}M_{G}.$ (104)
@@
Step 3 (Slab-wise warm-up bias bound). Assume $\pi^{\mathrm{warm}},\pi^{\star}\in\mathcal{U}$ and the slab-wise contraction $\|T(\pi_{1})-T(\pi_{2})\|_{k}\leq\rho(\tau_{k})\|\pi_{1}-\pi_{2}\|_{k}$ from Proposition 2. Let $\rho_{*}:=\max_{k}\rho(\tau_{k})<1$. Since $\pi_{\mathrm{proj}}=T(\pi^{\mathrm{warm}})$ and $\pi^{\star}=T(\pi^{\star})$, for each slab $D_{k}$ we have
$\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{k}=\|T(\pi^{\mathrm{warm}})-T(\pi^{\star})\|_{k}\leq\rho(\tau_{k})\,\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}\leq\rho_{*}\,\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}.$ (105)
@@
Step 4 (Residual identity and slab-wise closure). Define the warm-up aggregated stationarity residual (mixed-moment) on $D$ by
$r_{\mathrm{FOC,mix}}^{\mathrm{warm}}(z):=A_{\mathrm{warm}}(z)\,\pi^{\mathrm{warm}}(z)+G_{\mathrm{warm}}(z),\hskip 18.49988pt\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}:=\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{L^{2}(\mu)}.$
Also define the slab-wise residual sizes
$\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}:=\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{k},\hskip 18.49988pt\text{so that }\ (\varepsilon_{\mathrm{warm}}^{\mathrm{mix}})^{2}=\sum_{k=1}^{K}(\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}})^{2}.$
By construction of $\pi_{\mathrm{proj}}$,
$\pi^{\mathrm{warm}}-\pi_{\mathrm{proj}}=A_{\mathrm{warm}}^{-1}r_{\mathrm{FOC,mix}}^{\mathrm{warm}},$
hence on each slab
$\|\pi^{\mathrm{warm}}-\pi_{\mathrm{proj}}\|_{k}\leq\kappa\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (106)
@@
Now combine the triangle inequality on each slab with (105):
$\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}\leq\|\pi^{\mathrm{warm}}-\pi_{\mathrm{proj}}\|_{k}+\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{k}\leq\kappa\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}+\rho_{*}\,\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}.$
Since $\rho_{*}<1$, we close slab-wise:
$\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}\leq\frac{\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (107)
Plugging into (105) gives
@@
$\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{k}\leq\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (108)
Summing over slabs yields the global bias bound
@@
$\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{L^{2}(\mu)}\leq\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}.$ (109)
Step 5 (Finish). Combine (101), (104), and (109) to obtain

### v2

@@ D.1 Stability of the projection map $(A,G)\mapsto-A^{-1}G$
$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\;\leq\;2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}\;+\;2\kappa^{2}\bigl(M+\|\widetilde{G}\|_{L^{\infty}(D)}\bigr)\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$ (95)
@@
Throughout, $\|\cdot\|$ denotes the operator norm induced by the Euclidean norm.
Step 1: invertibility and uniform bound for $\widetilde{A}^{-1}$. For $\mu$-a.e. $z$, write
$\widetilde{A}(z)\;=\;A(z)\Bigl(I+A(z)^{-1}\bigl(\widetilde{A}(z)-A(z)\bigr)\Bigr).$
By (i) and (iii),
$\bigl\|A^{-1}(z)\bigl(\widetilde{A}(z)-A(z)\bigr)\bigr\|\;\leq\;\|A^{-1}\|_{L^{\infty}(D)}\,\|\widetilde{A}-A\|_{L^{\infty}(D)}\;\leq\;\kappa\cdot(2\kappa)^{-1}\;=\;\tfrac{1}{2}.$
Hence $I+A^{-1}(\widetilde{A}-A)$ is invertible and $\bigl\|\bigl(I+A^{-1}(\widetilde{A}-A)\bigr)^{-1}\bigr\|\leq(1-\tfrac{1}{2})^{-1}=2$. Therefore $\widetilde{A}(z)$ is invertible and
$\|\widetilde{A}^{-1}(z)\|\;=\;\bigl\|\bigl(I+A^{-1}(\widetilde{A}-A)\bigr)^{-1}A^{-1}(z)\bigr\|\;\leq\;2\|A^{-1}\|_{L^{\infty}(D)}\;\leq\;2\kappa,$
so $\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\leq 2\kappa$.
Step 2: difference of inverses. Using the identity $\widetilde{A}^{-1}-A^{-1}=\widetilde{A}^{-1}(A-\widetilde{A})A^{-1}$ and Hölder ($L^{\infty}\times L^{2}\times L^{\infty}\to L^{2}$),
$\|\widetilde{A}^{-1}-A^{-1}\|_{L^{2}(\mu)}\;\leq\;\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}\,\|A^{-1}\|_{L^{\infty}(D)}\;\leq\;2\kappa^{2}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$
Step 3: control error bound. Since $\pi=-A^{-1}G$ and $\widetilde{\pi}=-\widetilde{A}^{-1}\widetilde{G}$,
@@
Taking $L^{2}(\mu)$ norms and applying Hölder ($L^{\infty}\times L^{2}\to L^{2}$) yields
@@
Using Step 1, Step 2, and (ii),
$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\leq 2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}+2\kappa^{2}\,M\,\|\widetilde{A}-A\|_{L^{2}(\mu)}\leq 2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}+2\kappa^{2}\bigl(M+\|\widetilde{G}\|_{L^{\infty}(D)}\bigr)\|\widetilde{A}-A\|_{L^{2}(\mu)},$
which is (95). ∎
D.2 Slab-wise small-gain for the $q$-aggregated projection inputs
@@
Assume the working domain carries a time coordinate and, for concreteness, $D\subset[0,T]\times S$ and $\mu(dt,d\xi)=dt\otimes\nu(d\xi)$ for some reference measure $\nu$ on $S$. Fix a partition $0=t_{0}<t_{1}<\cdots<t_{K}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$ and define
$D_{k}:=D\cap([t_{k-1},t_{k}]\times S),\hskip 18.49988pt\mu_{k}:=\mu|_{D_{k}},\hskip 18.49988pt\|f\|_{k}:=\|f\|_{L^{2}(\mu_{k})}.$
@@
Let $U$ be a neighborhood of a locally optimal deployable $\theta$-blind policy $\pi^{\star}_{\mathrm{blind}}$ in $L^{2}(\mu)$ such that for all $\pi\in U$,
$\|A_{\pi}^{-1}\|_{L^{\infty}(D)}\leq\kappa,\hskip 18.49988pt\|G^{\mathrm{mix}}_{\pi}\|_{L^{\infty}(D)}\leq M_{G}$
for some constants $\kappa,M_{G}>0$. Assume that on each slab $D_{k}$ the $q$-aggregated projection inputs satisfy
$\|A_{\pi_{1}}-A_{\pi_{2}}\|_{k}\leq\overline{L}_{A}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988pt\|G^{\mathrm{mix}}_{\pi_{1}}-G^{\mathrm{mix}}_{\pi_{2}}\|_{k}\leq\overline{L}_{G}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},$ (96)
for all $\pi_{1},\pi_{2}\in U$, with constants $\overline{L}_{A},\overline{L}_{G}>0$ depending only on band data. Define the (population) projection map $T(\pi):=-A_{\pi}^{-1}G^{\mathrm{mix}}_{\pi}$ and
$\rho(\tau):=\Bigl(\kappa\,\overline{L}_{G}+\kappa^{2}M_{G}\,\overline{L}_{A}\Bigr)\tau^{1/2}.$
Then for each slab $D_{k}$ and all $\pi_{1},\pi_{2}\in U$,
$\|T(\pi_{1})-T(\pi_{2})\|_{k}\leq\rho(\tau_{k})\,\|\pi_{1}-\pi_{2}\|_{k}.$
In particular, if the partition is chosen so that
$\rho^{\ast}:=\max_{1\leq k\leq K}\rho(\tau_{k})<1,$
then $T$ is a strict contraction on every slab with constant at most $\rho^{\ast}$.
@@
Fix $k$ and $\pi_{1},\pi_{2}\in U$. Set $A_{i}:=A_{\pi_{i}}$ and $G_{i}:=G^{\mathrm{mix}}_{\pi_{i}}$. Then
$T(\pi_{1})-T(\pi_{2})=-A_{1}^{-1}(G_{1}-G_{2})-(A_{1}^{-1}-A_{2}^{-1})G_{2},\hskip 18.49988ptA_{1}^{-1}-A_{2}^{-1}=A_{1}^{-1}(A_{2}-A_{1})A_{2}^{-1}.$
Using Hölder ($L^{\infty}\times L^{2}\to L^{2}$) on $D_{k}$ together with $\|A_{i}^{-1}\|_{L^{\infty}(D)}\leq\kappa$ and $\|G_{2}\|_{L^{\infty}(D)}\leq M_{G}$ yields
@@
Applying (96) gives the claim with $\rho(\tau_{k})$. ∎
Remark 3 (Where the $\tau^{1/2}$ gain comes from; relation to huh2025a).
The short-time factor $\tau^{1/2}$ in (96) is the same parabolic smoothing mechanism used in our earlier slab-wise PGDPO analysis:
  • One represents the relevant value-gap / adjoint differences on a short slab via a Duhamel (semigroup) formula, differentiates in space, and applies Young-type convolution bounds.
  • In mixed norms, the gradient smoothing bound contributes a factor $\tau^{1/2-1/q}$ and the trivial embedding $L^{\infty}_{t}\hookrightarrow L^{q}_{t}$ contributes $\tau^{1/q}$, yielding net $\tau^{1/2}$.
A full PDE derivation of this short-time contraction philosophy is given in Huh et al. (2025b, Appendix B); see also the companion supplementary note to Huh et al. (2025b) (“On Proof of Theorem 4 (Policy Gap Bound)”) for a step-by-step slab contraction argument in a closely related setting. In the present paper the same smoothing principle is applied not to the costate update map directly, but to the induced $q$-aggregated projection inputs $(A_{\pi},G^{\mathrm{mix}}_{\pi})$.
D.3 Proof of Theorem 3
Proof of Theorem 3.
We keep the notation of the main text. Let $\pi^{\mathrm{warm}}$ be the warm-up (deployable, $\theta$-blind) policy and define the population $q$-aggregated projection inputs $(A_{\mathrm{warm}},G_{\mathrm{warm}})$ by
$A_{\mathrm{warm}}:=A_{\pi^{\mathrm{warm}}},\hskip 18.49988ptG_{\mathrm{warm}}:=G^{\mathrm{mix}}_{\pi^{\mathrm{warm}}}.$
Define the population projected policy
$\pi^{\mathrm{proj}}:=T(\pi^{\mathrm{warm}})=-A_{\mathrm{warm}}^{-1}G_{\mathrm{warm}}.$
@@ D.3 Proof of Theorem 3
Let $\widehat{A}$ and $\widehat{G}$ be the stage-2 (BPTT/Monte Carlo) estimators of $A_{\mathrm{warm}}$ and $G_{\mathrm{warm}}$ used to form the deployable stage-2 control
$\widehat{\pi}_{\mathrm{agg,mix}}:=-\widehat{A}^{-1}\widehat{G}$.
@@
Let $\pi^{\star}_{\mathrm{blind}}$ denote a locally optimal interior deployable $\theta$-blind policy, so in particular it satisfies the fixed-point (stationarity) identity
$\pi^{\star}_{\mathrm{blind}}=T(\pi^{\star}_{\mathrm{blind}}).$
Step 1 (Estimator-to-population projection stability). Assume we are in the stable inversion regime (as enforced in practice by ridge/diagnostics) so that the hypotheses of Proposition 1 apply on $D$ with $A=A_{\mathrm{warm}}$ and $G=G_{\mathrm{warm}}$. Then there exists a constant $C_{2}>0$ (depending only on the uniform inverse bound $\kappa$ and uniform $L^{\infty}$ bounds on the relevant $G$ terms) such that
$\|\widehat{\pi}_{\mathrm{agg,mix}}-\pi^{\mathrm{proj}}\|_{L^{2}(\mu)}\leq C_{2}\,\delta_{\mathrm{BPTT}},$ (97)
where one convenient choice is
$\delta_{\mathrm{BPTT}}:=\|\widehat{A}-A_{\mathrm{warm}}\|_{L^{2}(\mu)}+\|\widehat{G}-G_{\mathrm{warm}}\|_{L^{2}(\mu)}.$
[…]
@@
Step 2 (Slab-wise contraction for the population projection map). Assume $\pi^{\mathrm{warm}},\pi^{\star}_{\mathrm{blind}}\in U$ and let $\rho^{\ast}<1$ be the slab contraction constant of Proposition 2. Since $\pi^{\mathrm{proj}}=T(\pi^{\mathrm{warm}})$ and $\pi^{\star}_{\mathrm{blind}}=T(\pi^{\star}_{\mathrm{blind}})$, for each slab $D_{k}$ we have
$\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{k}=\|T(\pi^{\mathrm{warm}})-T(\pi^{\star}_{\mathrm{blind}})\|_{k}\leq\rho(\tau_{k})\,\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\rho^{\ast}\,\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}.$ (98)
@@
Step 3 (Residual identity and slab-wise warm-up deviation). Define the population mixed-moment aggregated stationarity residual of the warm-up policy
$r^{\mathrm{warm}}_{\mathrm{FOC,mix}}:=A_{\mathrm{warm}}\pi^{\mathrm{warm}}+G_{\mathrm{warm}},\hskip 18.49988pt\varepsilon^{\mathrm{mix}}_{\mathrm{warm}}:=\|r^{\mathrm{warm}}_{\mathrm{FOC,mix}}\|_{L^{2}(\mu)},\hskip 18.49988pt\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}:=\|r^{\mathrm{warm}}_{\mathrm{FOC,mix}}\|_{k}.$
By definition of $\pi^{\mathrm{proj}}$,
$\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}=\pi^{\mathrm{warm}}-(-A_{\mathrm{warm}}^{-1}G_{\mathrm{warm}})=A_{\mathrm{warm}}^{-1}\bigl(A_{\mathrm{warm}}\pi^{\mathrm{warm}}+G_{\mathrm{warm}}\bigr)=A_{\mathrm{warm}}^{-1}r^{\mathrm{warm}}_{\mathrm{FOC,mix}}.$
Hence, on each slab,
$\|\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}\|_{k}\leq\|A_{\mathrm{warm}}^{-1}\|_{L^{\infty}(D)}\,\|r^{\mathrm{warm}}_{\mathrm{FOC,mix}}\|_{k}\leq\kappa\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}.$ (99)
@@
Step 4 (Slab-wise closure). Combine the triangle inequality with (98) and (99):
$\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\|\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}\|_{k}+\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\kappa\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}+\rho^{\ast}\,\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}.$
Since $\rho^{\ast}<1$, we can close slab-wise:
$\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\frac{\kappa}{1-\rho^{\ast}}\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}.$ (100)
Plugging (100) into (98) yields
@@
$\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\frac{\rho^{\ast}\kappa}{1-\rho^{\ast}}\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}.$ (101)
Step 5 (Global bound and completion). Sum over slabs using $\|f\|_{L^{2}(\mu)}^{2}=\sum_{k}\|f\|_{k}^{2}$:
@@
$\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{L^{2}(\mu)}\leq\frac{\rho^{\ast}\kappa}{1-\rho^{\ast}}\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm}}.$
Finally, combine with (97) and the triangle inequality: